Delia can finish a job in 8 hours. Daisy can doit in 5 hours. if delia worked for 3 hours and then daisy was asked to help her finish it,? how long will daisy have to work with delia to finish the job?
Answer by mananth(16946)

Delia 8 hours
She does 1/ 8 job in 1 hour
Daisy 5 hours
she does 1/ 5 of thejob in 1 hour

Together they will do 1/ 8 + 1/ 5 in 1 hour
Together they will do 13/40 of the job in one hour
Delia works for 3 hours alone
so in 3 hours she does 3* 1/8 = 3/8 of the job.
Balance job = 1-3/8= 5/8 of the job
Together they do 13/40 of the job in 1 hour
They will take x hours to do 5/8 of the job
They will take (5/8) / (13/40) hours
=5/8 * 40/13
=25/13 hours
=
daisy will have to work with delia 25/13 hours